ϕ(λ) = \frac{1}{2πi} ∫_{c−i∞}^{c+i∞} exp (u ln u + λu) du  for c ≥ 0  (1)
λ = \frac{ϵ − ϵ̄}{ξ} − γ′ − β^{2} − ln \frac{ξ}{E_{max}}  (2)
γ = 0.577215 … (Euler′s constant)  (3)
γ′ = 0.422784 … = 1 − γ  (4)
ϵ, ϵ̄ = actual/average energy loss  (5)

Since 2 and 6 hold for arbitrary δc-vectors, it is clear that N (A) = R(B) and that when y = B(x) one has... ...the Pythagorians knew infinitely many solutions in integers to a^{2} + b^{2} = c^{2}. That no non-trivial integer solutions exist for a^{n} + b^{n} = c^{n} with integers n > 2 has long been suspected (Fermat, c.1637). Only during the current decade has this been proved (Wiles, 1995).

V π^{sr} = ⟨∑_{i} M_{i}V_{i}V_{i} + ∑_{i} ∑_{j>i} R_{ij}F_{ij}⟩  (6)

= ⟨∑_{i} M_{i}V_{i}V_{i} + ∑_{i} ∑_{j>i} ∑_{α} ∑_{β} r_{iαjβ}f_{iαjβ} − ∑_{i} ∑_{α} p_{iα}f_{iα}⟩